The coefficients of static and kinetic friction between a 389 N crate and the warehouse floor are 0.515 and 0.370, respectively. A worker gradually increases his horizontal push against this crate until it just begins to move and from then on maintains that same maximum push.

what is the acceleration of the crate after it has begun to move?
Answer #1

0.51 5 (189)는之00-33 N (300-33- 0-39 x)
Answer #2

SOLUTION :


Crate weight = 389 N 

Frictional force in static state = µ * 389 N = 0.515 * 389 = 200.335 N


Push (horizontal) needed to just start the movement 

= Frictional force in static state

= 200 .335 N 


This push force is continued even when the crate  moves .


So,


Net force acting on the crate in dynamic state

= Push force applied - Kinetic frictional force in moving state

= 200.335 - 0.370 * 389 

= 56.405 N


Now, 


m * a = Net force acting on the crate while moving


=> (389 / 9.8) * a = 56.405 

=> a = 56.405 * 9.8 / 389 = 1.421 m/s^2 

=> a = acceleration of crate movement = 1.421 m/s^2 (ANSWER)

answered by: Tulsiram Garg